Hello I need help in solving this problem. Thank you in advance and God Bless.

Solve the following problems.

1. The x-coordinates of an object at time t = 1.00 s and t = 4.00 s are 3.00 m and 5.00 m respectively. Calculate the average velocity of the object on the time interval t = 1.00 s to 4.00 s. Construct and interpret the position-time graph, velocity-time graph and acceleration-time graph.


1
Expert's answer
2021-01-14T10:37:40-0500

x co-ordinate at "t_1=1s" is "x_1 = 3m"


x co-ordinate at "t_2 =4s" is "x_2 = 5m"


Average velocity of the object is, "v_{av} = \frac{x_2-x_1}{t_2-t_1} = \frac{5-3}{4-1} = \frac{2}{3}m/s"


For the position-time graph, the curve will be linear passing through the points "(x_1,t_1), (x_2,t_2)" .

For the velocity-time graph, the curve will be parallel to the time axis as velocity is constant.

For the acceleration time graph, the curve will be along time axis as the acceleration of the object is zero.